OpenStudy (anonymous):

(again)Solve by Laplace transform: \[\frac{d^2y}{dx^2} + a \frac{dy}{dx} - 2a^2y=0\] y(0)=6, y'(0) = 0

OpenStudy (anonymous):

\[\frac{d^2y}{dx^2} + a \frac{dy}{dx} - 2a^2y=0\]\[sY'-0 +a(sY-6)-2a^2Y=0\]\[s(sY-6) +a(sY-6)-2a^2Y=0\]\[(s^2+as-2a^2)Y=6a+6s\]\[Y=\frac{6a+6s}{(s+2a)(s-a)}\]So far so good? Or...??

OpenStudy (experimentx):